# Deep Fourier-embedded Network for RGB and Thermal Salient Object Detection

### v2

@@ Abstract
The rapid development of deep learning has significantly improved salient object detection (SOD) combining both RGB and thermal (RGB-T) images. However, existing deep learning-based RGB-T SOD models suffer from two major limitations. First, Transformer-based models with quadratic complexity are computationally expensive and memory-intensive, limiting their application in high-resolution bi-modal feature fusion. Second, even when these models converge to an optimal solution, there remains a frequency gap between the prediction and ground-truth. To overcome these limitations, we propose a purely Fourier transform-based model, namely Deep Fourier-Embedded Network (DFENet), for accurate RGB-T SOD. To address the computational complexity when dealing with high-resolution images, we leverage the efficiency of fast Fourier transform with linear complexity to design three key components: (1) the Modal-coordinated Perception Attention, which fuses RGB and thermal modalities with enhanced multi-dimensional representation; (2) the Frequency-decomposed Edge-aware Block, which clarifies object edges by deeply decomposing and enhancing frequency components of low-level features; and (3) the Fourier Residual Channel Attention Block, which prioritizes high-frequency information while aligning channel-wise global relationships. To mitigate the frequency gap, we propose Co-focus Frequency Loss, which dynamically weights hard frequencies during edge frequency reconstruction by cross-referencing bi-modal edge information in the Fourier domain. Extensive experiments on four RGB-T SOD benchmark datasets demonstrate that DFENet outperforms fifteen existing state-of-the-art RGB-T SOD models. Comprehensive ablation studies further validate the value and effectiveness of our newly proposed components. The code is available at https://github.com/JoshuaLPF/DFENet.
Index Terms:
Fast Fourier transform, global dependency, RGB and thermal modalities, salient object detection.
@@ I Introduction
Salient object detection (SOD) identifies the most attention-grabbing objects in a scene, with broad applications spanning image compression [1], object recognition [2], scene perception [3], and more. Despite significant progress being made in RGB-based salient object detection [4, 5, 6], it remains challenging to separate salient objects from complex backgrounds or low-contrast scenes. To address this challenge, depth information was introduced to complement RGB features, enhancing the understanding of scenes [7, 8, 9, 10].
@@
However, the accuracy and stability of depth information are susceptible to environmental factors, such as varying light intensity, occlusions, and other conditions, limiting its applicability. In contrast, thermal infrared information offers distinct advantages in environmental adaptability. Even under adverse weather or low-light conditions, thermal images can effectively characterize objects by capturing their infrared radiations. As a result, RGB and thermal (RGB-T) SOD has gained increasing attention [11, 12, 13, 14].
@@
The Transformer architecture[15] has demonstrated excellent ability to acquire global dependencies, outperforming convolutional neural networks (CNNs) in various computer vision tasks [16, 17]. This success has led to the development of Transformer-based models for RGB-T SOD, which have achieved impressive performance [18, 19, 20, 21, 22, 23]. However, a significant challenge arises from the high memory consumption and computational complexity of Transformers, particularly when handling high-resolution bi-modal features and inputs [23]. To mitigate these computational expenses during training, existing Transformer-based models usually employ compromise strategies, such as constructing CNN-based fusion modules [18, 19, 21, 20], utilizing CNN-based backbones [21, 22], restricting Multi-head Self-attention (MHSA)-based modules to low-resolution feature fusion [7, 24, 25], and developing lightweight student models [26]. Despite these efforts, these approaches only partially explore global relationships, leading to less effective performance. This raises the question: is there a more efficient operation that can capture global relationships without compromising computational efficiency.
@@
Recently, benefiting from its powerful global representation capabilities, fast Fourier transform (FFT) [27] has demonstrated remarkable performance in image enhancement tasks [28, 29, 30, 31]. Notably, FFT is a global-based transform that represent all previous elements with each transformed element, making it memory-efficient. Inspired by this observation, we propose the purely FFT-based DFENet that aims to achieve high-resolution RGB-T SOD and bi-modal feature fusion by aligning global dependencies at every stage.
@@
The differences in imaging techniques between RGB and thermal modalities result in vast disparities in information distribution. As a result, while both modalities depict the same salient objects, they differ in the degree of saliency in the spatial dimension and exhibit distinct patterns in the channel dimension. To address this problem, we propose Modal-coordinated Perception Attention (MPA), a Transformer-like module designed to align and fuse the two modalities effectively across both spatial and channel dimensions through a re-embedding strategy.
@@
Furthermore, neural networks inherently tend to lose vital visual details as they deepen, while reconstructing high-frequency edge information is prone to spectral bias [32], causing models to prioritize learning easier frequencies over the hard ones. However, existing edge modules based on spatial domain learning tend to generate sparse spectra [19, 33, 26], treating all frequencies equally and making it difficult to accurately characterize edges. To address these issues, inspired by [34, 35, 36], we propose the Co-focus Frequency Loss (CFL) to guide FEB in the frequency domain while optimizing the entire model. In CFL, we measure the prediction and ground-truth using the Fourier frequency distance and introduce original edge frequency information of RGB and thermal modalities as a cross-reference to dynamically weight hard frequencies. This approach leads to refined edge features that further improve the quality of the saliency map in spatial domain learning.
By integrating these components, DFENet efficiently fuses RGB and thermal modalities by utilizing complementary relationships from the frequency perspective, ultimately producing saliency maps with high accuracy.
In summary, the main contributions of this paper are as follows:
  • We propose DFENet, an FFT-based model designed to achieve high-resolution RGB-T SOD and bi-modal feature fusion while minimizing memory consumption. Through comprehensive ablation studies, we verify the effectiveness of the proposed components, including MPA, FEB, FRCAB, and CFL.
  • We conduct extensive experiments on four benchmark datasets and show that our proposed DFENet achieves superior performance and robustness compared to fifteen existing state-of-the-art models. Notably, DFENet is able to process bi-modal inputs with resolution up to $512\times 512$ on an NVIDIA Tesla P100 GPU with 16GB of memory.
  • To the best of our knowledge, DFENet is the first FFT-based supervised model for SOD tasks. This work demonstrates the strong potential of FFT-based operations for dense prediction, highlighting their unique advantages while also exploring their limited compatibility with spatial domain operations.
@@ II-A RGB and Depth Salient Object Detection (RGB-D SOD)
CNN-based RGB-D SOD models typically leverage foundational attention mechanisms, such as convolutional block attention module [37] utilized in [38, 39, 40, 41, 42, 43], dilated convolution [44] employed in [45, 46] and atrous spatial pyramid pooling [47] utilized in [48, 49, 50]. In addition, Ji et al. [9] proposed an unsupervised network that automatically generates pseudo-labels through iterative refinement. In [8], a depth calibration and fusion framework was proposed to mitigate the effect of noise and ambiguity in depth images. Furthermore, Yao et al. [51] augmented semantic representation by integrating depth maps into the single-stream encoder.
To address the limitations imposed by the local connectivity of CNNs, several Transformer-based models [15, 17] have been proposed. Liu et al. [52] introduced three Vision Transformers [16] with shared weights to refine high-level fused features. The pure Transformer networks were proposed in [7, 25] for simultaneous salient object and edge detection. Sun et al. [24] employed the Swin Transformer [17] as a dual-stream encoder, utilizing Multi-head Self-attention (MHSA) to handle high-level features. Wu et al. [10] proposed a cross-modality interaction parallel Transformer to fuse RGB and depth features, complemented by a contrast learning strategy. In contrast to these Transformer-based spatial domain models, our proposed DFENet employed Fourier Transform (FFT)-based operations to solve bi-modal SOD in the frequency domain.
@@ II-B RGB and Thermal Salient Object Detection (RGB-T SOD)
In complex environments with varying lighting conditions or occlusions, depth information can become unreliable. To address this limitation, researchers [11, 12, 13] have introduced thermal modality as an alternative to depth information for SOD. Thermal imaging captures the infrared radiation emitted by objects and is more adaptable to different environmental conditions.
Similar to RGB-D SOD (Sec. II-A), recent efforts in RGB-T SOD primarily focus on inter-modal fusion. Several CNN-based fusion strategies have been proposed [53, 54, 55, 56, 14, 57, 58, 59] to combine multi-modal, multi-scale, and multi-granular features. For example, Song et al. [14] utilized graph structures to reduce the impact of negative information. Inspired by style transfer, Chen et al. [57] devised a modal transfer mechanism to bridge complementary features. In addition, Tu et al. [60] proposed a weak alignment model for RGB and thermal modalities. Cong et al. [61] re-examined the role of thermal modality in relation to light variations.
To capture long-range relationships, researchers have introduced Transformers for various purposes, including use as backbones [18, 19, 20], post-processing of fused features [21, 20], or feature fusion [22]. In [22], Pang et al. proposed a cross-modal and cross-scale Transformer with a token re-embedding strategy to alleviate complexity. However, the substantial memory consumption of MHSA makes it challenging to apply Transformers in each stage of the model. A pure transformer model [23] using the Swin Transformer [17] in both encoder and decoder required four RTX3090 GPUs for training due to its computational demands.
To address the expensive computation, several studies [62, 63, 33] have explored lightweight modeling, but their performance has often fallen short of expectations. Zhou et al. [26] employed knowledge distillation, designing a wavelet transform-based WaveNet to acquire semantic and geometric knowledge from a complex teacher model [19]. Despite its merits, the wavelet transform is inherently locally based and fails to capture long-range dependencies effectively.
Considering these limitations, we design a purely FFT-based DFENet. The FFT is a globally based transform that has an advantage over MHSA and the wavelet transform in efficiently obtaining global relationships. This property allows each component of DFENet to access and optimize global relationships effectively. Moreover, DFENet offers memory efficiency and supports input resolution up to $512^{2}$ on an NVIDIA Tesla P100 GPU with 16GB of memory (see Sec. IV-C for details).
@@ II-C Fast Fourier Transform (FFT)-based Networks
Despite the recent advancement achieved by MHSA-based Transformers [16, 17] in various tasks, the trade-off between the perceived field size and computational cost remains a challenge. To address this problem, researchers have explored alternative approaches based on the FFT. For example, Rao et al. [64] proposed an FFT-based global filter to learn long-distance dependencies with linear complexity. Guibas et al. [65] constructed a token mixer as a globally continuous convolution, while Tatsunami et al. [66] incorporated the concept of data correlation into the global filter to dynamically extract image features. Despite these advancements, existing works have primarily focused on the spatial dimension, overlooking the exploration of intrinsic discriminative features in the channel dimension. Zhou et al. [67] addressed this limitation by designing two branches to perform FFT in both spatial and channel dimensions for image restoration.
Meanwhile, researchers have also applied FFT-based models for various downstream tasks, including image enhancement [28, 29, 30, 31], image inpainting [68], image compression [36], and image reconstruction [69, 70]. However, there is a notable scarcity of research on dense prediction tasks. Existing dense prediction models [71, 72] typically focus on high- and low-frequency components of the Fourier signal, neglecting deeper components such as amplitude and phase.
In this paper, we proposed DFENet to develop FFT for addressing dense prediction tasks, specifically SOD. To the best of our knowledge, this is the first FFT-based supervised model specifically designed for this task.
Figure 1: The framework of our proposed DFENet. RGB and thermal images a’re first extracted initial features $\left\{{{r_{i}}}\right\}_{i=1}^{4}$ and $\left\{{{t_{i}}}\right\}_{i=1}^{4}$ by a dual-stream encoder. Next, the bi-modal features enter the Modal-coordinated Perception Attention (MPA) to bridge the complementary information and fuse. During decoding, the fused features $\left\{{{f_{i}}}\right\}_{i=1}^{4}$ are integrated in a low- to high-resolution manner by Fourier Residual Channel Attention Block (FRCAB) to obtain decoder features $\left\{{{d_{i}}}\right\}_{i=1}^{4}$. Meanwhile, the shallow features $\left\{{{t_{i},f_{i}}}\right\}_{i=1}^{2}$ are fed into the Frequency-decomposed Edge-aware Block (FEB) to obtain elaborate edge features $\left\{{{e_{i}}}\right\}_{i=1}^{3}$ for guiding the decoding process in generating the accurate saliency map $\mathcal{S}$. In the training stage, the Co-focus Frequency Loss (CFL) ${{\cal L}_{CFL}}$ favoring the synthesis of difficult frequencies and spatial-domain losses are employed together to supervise the generation of the high-quality edge map $\mathcal{E}$.
@@ III-A Overview
In this section, we introduce the proposed DFENet, a novel architecture designed for RGB and thermal salient object detection (RGB-T SOD), as illustrated in Fig. 1.
The network begins with a dual-stream encoder based on fast Fourier transform (FFT)-based CDFFormer-m [66], which processes the RGB and thermal modalities separately to extract set of features at four different scales, denoted as $\left\{{{r_{i}}}\right\}_{i=1}^{4}$ and $\left\{{{t_{i}}}\right\}_{i=1}^{4}$, respectively.
The extracted features are then processed by our Modal-coordinated Perception Attention (MPA) (Sec. III-B), which is designed to obtain cross-modal fusion features $\left\{{{f_{i}}}\right\}_{i=1}^{4}$.
Subsequently, $\left\{{{t_{i}},{f_{i}}}\right\}_{i=1}^{2}$ are input into our Frequency-decomposed Edge-aware Block (FEB) (Sec. III-C), which is designed to extract detailed edge features $\left\{{{e_{i}}}\right\}_{i=1}^{3}$.
In the decoding stage, we employ a pyramid design to handle object scale variations. Each layer of the decoder concatenates the corresponding fusion feature ${f_{i}}$ with features from all previous layers. These concatenated features are then refined using our Fourier Residual Channel Attention Block (FRCAB) (Sec. III-D), resulting in refined features $\left\{{{d_{i}}}\right\}_{i=1}^{4}$.
Finally, the network is optimized using a bi-domain learning approach (Sec. III-E), which combines the proposed Co-focus Frequency Loss (CFL) ${{\cal L}_{CFL}}$, binary cross-entropy loss (BCE) [73] ${{\cal L}_{BCE}}$, and intersection-over union loss (IoU) [74] ${{\cal L}_{IoU}}$. ${{\cal L}_{CFL}}$ in the frequency domain facilitates global optimization, prioritizing the resolution of hard frequencies during the edge-frequency reconstruction process. Meanwhile, ${{\cal L}_{BCE}}$ and ${{\cal L}_{IoU}}$ in the spatial domain achieve substantial pixel-wise saliency prediction $\mathcal{S}$.

### v3

@@ Abstract
The rapid development of deep learning has significantly improved salient object detection (SOD) combining both RGB and thermal (RGB-T) images. However, existing Transformer-based RGB-T SOD models with quadratic complexity are memory-intensive, limiting their application in high-resolution bimodal feature fusion. To overcome this limitation, we propose a purely Fourier Transform-based model, namely Deep Fourier-embedded Network (FreqSal), for accurate RGB-T SOD. Specifically, we leverage the efficiency of Fast Fourier Transform with linear complexity to design three key components: (1) To fuse RGB and thermal modalities, we propose Modal-coordinated Perception Attention, which aligns and enhances bimodal Fourier representation in multiple dimensions; (2) To clarify object edges and suppress noise, we design Frequency-decomposed Edge-aware Block, which deeply decomposes and filters Fourier components of low-level features; (3) To accurately decode features, we propose Fourier Residual Channel Attention Block, which prioritizes high-frequency information while aligning channel-wise global relationships. Additionally, even when converged, existing deep learning-based SOD models’ predictions still exhibit frequency gaps relative to ground-truth. To address this problem, we propose Co-focus Frequency Loss, which dynamically weights hard frequencies during edge frequency reconstruction by cross-referencing bimodal edge information in the Fourier domain. Extensive experiments on ten bimodal SOD benchmark datasets demonstrate that FreqSal outperforms twenty-nine existing state-of-the-art bimodal SOD models. Comprehensive ablation studies further validate the value and effectiveness of our newly proposed components. The code is available at https://github.com/JoshuaLPF/FreqSal.
@@ I Introduction
Salient object detection (SOD) identifies the most attention-grabbing objects in a scene, with broad applications spanning image compression [1], object recognition [2], scene perception [3], and more. Despite significant progress being made in RGB-based SOD [4, 5], it remains challenging to separate salient objects from complex backgrounds or low-contrast scenes. To address this challenge, depth information was introduced to complement RGB features, enhancing the understanding of scenes [6, 7, 8, 9, 10, 11].
@@
However, the accuracy and stability of depth information are susceptible to environmental factors, such as varying light intensity, occlusions, and other conditions, limiting its applicability. In contrast, thermal infrared information offers distinct advantages in environmental adaptability. Even under adverse weather or low-light conditions, thermal images can effectively characterize objects by capturing their infrared radiation. As a result, RGB and thermal (RGB-T) SOD has gained increasing attention [12, 13, 14, 15, 16].
@@
The Transformer architecture[17] has demonstrated excellent ability to acquire global dependencies, outperforming convolutional neural networks (CNNs), which are constrained by local receptive fields, in various computer vision tasks [18, 19]. This success has led to the development of Transformer-based models for RGB-T SOD, which have achieved impressive performance [20, 21, 22, 23, 24, 25]. However, a significant challenge arises from the high memory consumption of Transformers, particularly when handling high-resolution bimodal features [25, 26]. To mitigate the memory footprint for training with limited computational resources, existing Transformer-based accuracy-targeted models [^1] Accuracy-targeted models focus on achieving high performance and are typically heavyweight, whereas lightweight models emphasize efficiency. usually employ compromise strategies, such as sacrificing the exploration of relationships between higher resolution bimodal features [6, 25, 27, 28, 29], constructing CNN-based fusion modules [20, 21, 23, 22, 30], and utilizing CNN-based backbones [23, 24, 30] [^2] The simplistic structure diagrams of existing Transformer-based accuracy-targeted models are presented in the Appendix.. These compromise strategies result in models only partially explore global relationships, leading to less effective performance. This raises the question: Is there a more efficient operation that captures global relationships across each part of the model while avoiding excessive memory consumption?
@@
Recently, benefiting from its powerful global representation capabilities, Fast Fourier transform (FFT) [31] has shown remarkable performance in image enhancement tasks [32, 33, 34, 35]. Importantly, FFT is a global-based transform that represents all previous elements with each transformed element, making it memory-efficient.
Figure 1: Frequency spectrum visualization: (a) RGB and (b) thermal images, with (c) and (d) showing their corresponding frequency spectrograms.
Inspired by this observation, we propose the purely FFT-based architecture, FreqSal, aiming to align global dependencies at each stage—including encoding, modal fusion, decoding, and edge detection—while achieving high-resolution bimodal feature fusion and accurate RGB-T SOD.
@@
The differences in imaging techniques between RGB and thermal modalities result in vast disparities in information distribution. For example, the frequency spectrograms of RGB and thermal images in Fig. 1 reveal notable differences, reflecting their distinct frequency distributions of saliency. Specifically, while both modalities depict the same salient objects, they differ in the degree of saliency in the spatial dimension and exhibit distinct patterns in the channel dimension [21, 36]. To address this problem, we propose Modal-coordinated Perception Attention (MPA), a Transformer-like module that aligns and fuses the two modalities across both spatial and channel dimensions in the frequency domain through a re-embedding strategy.
@@
Furthermore, existing deep learning-based SOD models [27, 28, 29, 26, 37, 38] often lose vital visual details as the network deepens. In addition, reconstructing high-frequency edge information is prone to spectral bias [39], causing these models to favor learning easier frequencies over the hard ones. However, existing edge modules based on spatial domain learning tend to generate sparse spectra [21, 40, 41], treating all frequencies equally and making it difficult to accurately characterize edges. To address these issues, inspired by [42, 43, 44], we propose the Co-focus Frequency Loss (CFL) to guide FEB in the frequency domain while optimizing the entire model. In CFL, we measure the prediction and ground-truth using the Fourier frequency distance and introduce original edge frequency information of RGB and thermal modalities as a cross-reference to dynamically weight hard frequencies. This frequency domain learning approach leads to refined edge features that further improve the quality of the saliency map in spatial domain learning.
By integrating these components, FreqSal efficiently fuses RGB and thermal modalities by utilizing complementary relationships from the frequency perspective, ultimately producing saliency maps with high accuracy. In summary, the main contributions of this paper are as follows:
  • We propose FreqSal, an FFT-based model designed to achieve high-resolution bimodal feature fusion and RGB-T SOD while minimizing memory consumption. Through comprehensive ablation studies, we verify the effectiveness of the proposed components, including MPA, FEB, FRCAB, and CFL.
  • We conduct extensive experiments on the ten benchmark datasets and show that our proposed FreqSal achieves superior performance and robustness over twenty-nine existing state-of-the-art models. Notably, FreqSal is able to process bimodal inputs with resolution up to $512\times 512$ on an NVIDIA Tesla P100 GPU with 16GB of memory.
  • To the best of our knowledge, FreqSal is the first FFT-based supervised model for SOD tasks. This work demonstrates the potential of FFT-based operations for dense prediction, highlighting their unique advantages while also exploring their limited compatibility with spatial domain operations.
@@ II-A RGB and Depth Salient Object Detection (RGB-D SOD)
CNN-based RGB-D SOD models typically leverage foundational attention mechanisms, such as convolutional block attention module [45] utilized in [46, 47], dilated convolution [48] employed in [49, 50] and atrous spatial pyramid pooling [51] utilized in [52, 53]. In addition, Ji et al. [9] utilized depth information to guide recurrent attention networks. Furthermore, a three-stream learning network was proposed in [10] to automatically filter out low-quality depth maps.
To address the limitations imposed by the local connectivity of CNNs, several Transformer-based models [17, 19] have been proposed. Liu et al. [54] introduced three Vision Transformers [18] with shared weights to refine high-level fused features. The pure Transformer networks were proposed in [6, 27] for simultaneous salient object and edge detection. Several works [55, 56, 57] employed the Swin Transformer [19] as a dual-stream encoder to obtain compact feature representations. Jin et al. [58] proposed a two-stage “simple-to-complex” strategy for learning RGB-D samples. In contrast to these Transformer-based spatial domain models, our proposed FreqSal employed FFT-based operations to solve bimodal SOD in the frequency domain.
@@ II-B RGB and Thermal Salient Object Detection (RGB-T SOD)
In complex environments with varying lighting conditions or occlusions, depth information become unreliable. To address this limitation, researchers [12, 13, 14] have introduced thermal modality as an alternative to depth information for SOD. Thermal imaging captures the infrared radiation emitted by objects and is more adaptable to different environmental conditions.
Similar to RGB-D SOD (Sec. II-A), recent efforts in RGB-T SOD primarily focus on inter-modal fusion. Several CNN-based fusion strategies have been proposed [59, 60, 15, 61, 62, 30] to combine multi-modal, multi-scale, and multi-granular features. In addition, Tu et al. [63] proposed a weak alignment model for RGB and thermal modalities. Cong et al. [64] re-examined the role of thermal modality in relation to light variations.
To capture long-range relationships, researchers have introduced Transformers for various purposes, including use as backbones [20, 21, 22, 26, 65, 66], post-processing of fused features [23, 22], or feature fusion [24]. In [67, 68], triple-flow networks were proposed to mitigate the effects of modal bias and inherent noise. Wang et al. [29] proposed a semantics-guided attention to identify the relevance of unaligned modalities. However, the substantial memory consumption of Multi-head Self-attention (MHSA) makes it challenging to apply Transformers in each stage of the model. A pure transformer model [25] using the Swin Transformer [19] in both encoder and decoder required four RTX3090 GPUs for training.
To address the expensive computation, several studies [69, 70, 40, 71, 72] have explored lightweight modeling, but their performance has often fallen short of expectations. Zhou et al. [41] designed a wavelet transform-based model to acquire saliency knowledge from a teacher model [21]. Despite its merits, the wavelet transform is inherently locally based and fails to capture global dependencies effectively.
Considering these limitations, we design a purely FFT-based FreqSal. The FFT is a global transform that has an advantage over MHSA and the wavelet transform in efficiently obtaining global relationships. This property allows each component of FreqSal to access and optimize global relationships effectively. Moreover, FreqSal offers memory efficiency and supports input resolution up to $512^{2}$ on an NVIDIA Tesla P100 GPU with 16GB of memory (see Sec. IV-E for details).
@@ II-C Fast Fourier Transform (FFT)-based Networks
Despite the recent advancement achieved by MHSA-based Transformers [18, 19] in various tasks, the trade-off between the perceived field size and computational cost remains a challenge. To address this problem, researchers have explored alternative approaches based on the FFT. For example, Rao et al. [73] proposed an FFT-based global filter to learn long-distance dependencies with linear complexity. Guibas et al. [74] constructed a token mixer as a globally continuous convolution, while Tatsunami et al. [75] incorporated the concept of data correlation into the filter to dynamically extract image features. In parallel, researchers have also applied FFT-based models for various downstream tasks, including image enhancement [32, 33, 34, 35], image inpainting [76], image compression [44], and image reconstruction [77, 78]. However, there is a notable scarcity of research on dense prediction tasks. Existing dense prediction models [79, 80] typically focus on high- and low-frequency components of the Fourier signal, neglecting deeper aspects such as amplitude and phase.
In this paper, we proposed FreqSal to develop FFT for addressing dense prediction tasks, specifically SOD. To the best of our knowledge, this is the first FFT-based supervised model specifically designed for this task.
Figure 2: The framework of our proposed FreqSal. RGB and thermal images are first extracted initial features $\left\{{{r_{i}}}\right\}_{i=1}^{4}$ and $\left\{{{t_{i}}}\right\}_{i=1}^{4}$ by a dual-stream encoder. Next, the bimodal features enter the Modal-coordinated Perception Attention (MPA) to bridge the complementary information and fuse. During decoding, the fused features $\left\{{{f_{i}}}\right\}_{i=1}^{4}$ are integrated in a low- to high-resolution manner by Fourier Residual Channel Attention Block (FRCAB) to obtain decoder features $\left\{{{d_{i}}}\right\}_{i=1}^{4}$. Meanwhile, the shallow features $\left\{{{t_{i},f_{i}}}\right\}_{i=1}^{2}$ are fed into the Frequency-decomposed Edge-aware Block (FEB) to obtain elaborate edge features $\left\{{{e_{i}}}\right\}_{i=1}^{3}$ for guiding the decoding process in generating the accurate saliency map $\mathcal{S}$. In the training stage, the Co-focus Frequency Loss (CFL) ${\cal L}_{CFL}$} favoring the synthesis of difficult frequencies and spatial-domain losses are employed together to supervise the generation of the high-quality edge map $\mathcal{E}$.
@@ III-A Overview
In this section, we introduce the proposed FreqSal, a novel architecture designed for RGB-T SOD, as illustrated in Fig. 2. The network begins with a dual-stream encoder based on FFT-based CDFFormer-M [75], which processes the RGB and thermal modalities separately to extract a set of features at four different scales, denoted as $\left\{{{r_{i}}}\right\}_{i=1}^{4}$ for RGB features and $\left\{{{t_{i}}}\right\}_{i=1}^{4}$ for thermal features. The extracted features are then processed by our MPA (Sec. III-B), which is designed to obtain cross-modal fusion features $\left\{{{f_{i}}}\right\}_{i=1}^{4}$. Subsequently, $\left\{{{t_{i}},{f_{i}}}\right\}_{i=1}^{2}$ are input into our FEB (Sec. III-C), which is designed to extract detailed edge features $\left\{{{e_{i}}}\right\}_{i=1}^{3}$. In the decoding stage, we employ a pyramid design to handle object scale variations. Each layer of the decoder concatenates the corresponding fusion feature $f_{i}$} with features from all previous layers. These concatenated features are then refined using our FRCAB (Sec. III-D), resulting in refined features $\left\{{{d_{i}}}\right\}_{i=1}^{4}$. Finally, the network is optimized using a bi-domain learning approach (Sec. III-E), which combines the proposed CFL ${\cal L}_{CFL}$}, binary cross-entropy loss (BCE) [81] ${\cal L}_{BCE}$}, and intersection-over union loss (IoU) [82] ${\cal L}_{IoU}$}. ${\cal L}_{CFL}$} in the frequency domain facilitates global optimization, prioritizing the resolution of hard frequencies during the edge-frequency reconstruction process. Additionally, ${\cal L}_{BCE}$} and ${\cal L}_{IoU}$} in the spatial domain achieve substantial pixel-wise saliency prediction $\mathcal{S}$.